A. 0.4869 B. 0.5904 C. 0.6234 D. 0.5834 Explanation:Advertisement P(winning prize atleast on one ticket) = 1 – P(“Losing on all tickets”) = 1 – (0.8)4 = (1 + (0.8)2)(1 – (0.8)2) = (1.64)(0.36) = 0.5904 Advertisement